OpenStudy (anonymous):

On a sunny day,the daily profit, P dollars,of a popcorn vendor is modelled by the equation P= -60(x-4)^2 +120 where x is the price of a bag of popcorn. a) What should the price be to maximize daily profit ? b) What is the maximum amount of daily profit ? On a sunny day,the daily profit, P dollars,of a popcorn vendor is modelled by the equation P= -60(x-4)^2 +120 where x is the price of a bag of popcorn. a) What should the price be to maximize daily profit ? b) What is the maximum amount of daily profit ? @Mathematics

OpenStudy (anonymous):

a) dp/dx = -120(x-4) 0 = -120x + 480 120x= 480 x = 4 b) p(4) = -60(4-4)^2 + 120 = 120

OpenStudy (anonymous):

how did you get that ? like what are the steps ?

OpenStudy (anonymous):

a) derivative of p .. p'(x) and equate it to zero .. to find minimum/maximum b) plug the value from a to get the maximum amount of daily profit ..